OpenStudy (anonymous):

PLEASE ANYONE!! Define the Laplace transform L{f(t)} = F (s) of the function f(t) L{df/dt}= sF(s)−f(0)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Isn't it just int 0 & inf e^-st(f(t) dt ?

OpenStudy (anonymous):

Yeah I also think its Lapalce: L\[L\left\{ f \right\}_{s}=\int\limits_{0}^{\infty}e ^{-st}f(t)dt\]

OpenStudy (dumbcow):

http://www.bing.com/videos/search?q=LaPlace+transform&mid=10B8A6DA9A7F295BD39410B8A6DA9A7F295BD394&view=detail&FORM=VIRE5 i dont know but here is some reference you can use